OpenStudy (anonymous):

How do I solve a Chinese Remainder question such as : Given x, x≡3 (mod 7) x≡4 (mod 5) x≡0 (mod 3) Find x.

OpenStudy (asnaseer):

There is an excellent video here that explains this theorem and how to apply it: http://www.youtube.com/watch?v=3PkxN_r9up8

OpenStudy (asnaseer):

Did that video help?

OpenStudy (anonymous):

U take the moduli pairwise: 7 and 5*3 = 15 5 and 7*3 = 21 3 and 7*5 = 35 Then u need to find combinations of (7,15) (5,21) and (3,35) that give 1 eg -2*7 + 1*15 = 1 giving a factor 15 Then multiply the 3 factors by 3,4,0 respectively (the zero makes life easier in this case) and add them to get a solution (mod 3*5*7)

OpenStudy (anonymous):

thanks @asnaseer and @estudier

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

np